# DiffOG: Differentiable Policy Trajectory Optimization with Generalizability

### v4

@@ I Introduction
Imitation learning [16] has emerged as a popular paradigm for endowing robots with complex manipulation skills by leveraging human demonstrations. By formulating a supervised learning problem that maps sensor observations directly to actions via neural networks, imitation learning-based visuomotor policies have demonstrated effectiveness across a broad spectrum of tasks [5, 6, 36, 8, 22]. Yet, these learned policies often exhibit notable limitations when applied to real-world settings, where safety, robustness, and strict constraint satisfaction are paramount. In contrast to classical model-based trajectory optimization methods which explicitly account for motion constraints and can ensure smooth and reliable trajectories [20], imitation learning approaches may produce suboptimal or jerky actions. Such trajectories can reduce the robustness of the whole robotic systems and raise critical safety concerns.
Meanwhile, various challenges in imitation learning, such as the representation of multi-modal action distributions [5, 10] and training large, generalized policies on extensive datasets [14, 15, 3], have been the focus of ongoing research. Despite these advancements, ensuring the generation of high-quality and constraint-compliant trajectories for neural network-based policies remains a significant and open challenge. Unlike traditional model-based motion generation methods, learning-based approaches tend to be less controlled and more difficult to interpret. For instance, learning-based policies lack mechanisms to directly control trajectory smoothness or impose motion constraints through tunable parameters.
@@
We evaluated DiffOG across 11 simulated tasks and 2 real-world tasks. The results demonstrate that DiffOG significantly enhances the trajectory quality of visuomotor policies with generalizability while having minimal impact on policy performance, outperforming trajectory processing baselines such as constraint clipping and penalty-based trajectory optimization. Furthermore, DiffOG achieves superior performance compared to existing constrained visuomotor policy [33].
@@ II-A Supervised Policy Learning
Supervised policy learning entails training a policy using supervised learning techniques on a dataset of pre-collected demonstrations [16]. Recent advancements in machine learning algorithms and data collection frameworks have substantially advanced the capabilities and effectiveness of supervised policy learning in manipulation applications. For example, various data collection frameworks and devices have been proposed in [6, 36, 8, 22, 21, 25], enabling efficient data collection and the training of effective manipulation policies. To better capture the multi-modal action distributions present in demonstration data, prior work have explored various techniques, such as using diffusion models [5], VQ-VAE [10], and energy-based models [7]. Furthermore, various effective action representations for supervised policy learning have been proposed, such as flow by point tracking [31, 28] and equivariant action representation [26]. Moreover, many work have contributed to training more general and scalable policies on large-scale datasets [14, 15, 3, 12].
@@ II-B Policies with Differentiable Optimization
The integration of differentiable optimization into policy learning has garnered significant attention due to its potential to combine model-based formulations with the representational power of neural networks. Previous studies have investigated the integration of model-based structures with policies through differentiable optimization, including differentiable model predictive control (MPC) [2] and the Koopman operator framework [19]. However, these methods are predominantly designed for low-dimensional observations, which limits their flexibility and applicability to the general formulation of visuomotor policies. The work in [32] presents a tactile-reactive grasping controller that integrates an image encoder with differentiable MPC. Although effective for specific tactile-based grasping tasks, this approach lacks the formulation required for broader policy learning frameworks. In the domain of obstacle avoidance, an end-to-end learning framework incorporating differentiable optimization is proposed in [30]. However, this approach is specifically designed for navigation tasks and lacks applicability to robotic manipulation scenarios and visuomotor policies.
@@
Riemannian motion policies [18] and their extensions [11, 4, 17] advance the generation of motions in high-dimensional spaces with complex, nonlinear dynamics into end-to-end robot learning by leveraging automatic differentiation and Riemannian geometry. However, in the context of policy learning, Riemannian motion policies are primarily designed for tasks involving low-dimensional observations, such as those demonstrated through kinesthetic teaching [17]. In contrast, DiffOG is tailored for a broader range of visuomotor policies, refining policies that utilize camera observations as input. DiffOG focuses on generating task-completing trajectories while enhancing smoothness and ensuring compliance with hard constraints.
@@
DiffTORI [24] integrates differentiable trajectory optimization into the policy representation to generate actions, with a focus on capturing multi-modal action distributions in the context of imitation learning. However, DiffTORI lacks explicit and rigorous formulations for constraints and smoothness, limiting its ability to optimize trajectory smoothness and enforce compliance with hard constraints. In contrast, DiffOG incorporates a rigorously designed optimization layer that not only enables actions to satisfy hard constraints, but also enhances the interpretability of the policy.
@@
Leto [33], by employing differentiable optimization, enables end-to-end training and inference that embedded trajectory optimization, allowing constraints to be incorporated as part of the training objectives. However, Leto was designed as a discriminative model, which limits its ability to represent multi-modal action distributions [7]. Its use of an optimization layer as the action head makes it difficult to integrate with commonly-used diffusion-based action heads [5, 14, 12, 35], significantly reducing its flexibility and applicability. In comparison, DiffOG offers greater flexibility and enhanced representational capacity, enabling it to achieve superior policy performance and adapt more effectively to a wider range of tasks.
@@ III-B Action Space
Actions fall into two categories: discrete actions, such as controlling grasping and releasing using discrete values [13], and continuous actions, such as the motion of the robot. Discrete actions, by their nature, cannot form a continuous trajectory and therefore do not require trajectory optimization. To address this distinction, we introduce a selection matrix $S\in\mathbb{R}^{D_{c}\times D_{a}}$ that identifies the action dimensions to be considered for trajectory optimization. Specifically, $S$ is used to select: 1) degrees of freedoms (DOFs) that are inherently continuous and capable of forming a trajectory, and 2) DOFs that require trajectory optimization in practical applications. The resulting filtered action variable is expressed as $c=Sa\in\mathbb{R}^{D_{c}}$. In this paper, we filter out the grasping action using $S$, as grasping typically does not require trajectory motion constraints and optimization.
Recently, considerable papers demonstrate that using the training objective of predicting a sequence of actions results in a more effective policy compared to predicting a single action [5, 36, 10]. Therefore, at time step $t$, the action sequence output of policies adhere to the format of
@@ III-C Differentiable Trajectory Optimization
In the optimization problem (6), $\mathbf{Q}$ is a weight matrix, $\alpha$ is a scalar, and $d_{\text{max}},d_{\text{min}}$ are constraint bounds for the derivatives of the robot actions. Intuitively, $\alpha$ reflects the degree to which the trajectory smoothness is emphasized. When the sum of the derivatives of the robot actions $\hat{\mathbf{c}}_{t}$ in the generated trajectory $\mathbf{y}_{t}$ is large, it indicates that the trajectory is less smooth. For instance, excessive acceleration or velocity contributes to an increase in the cost function. Therefore, we call $\alpha$ as smoothing weight. Optimization problem (6) aims to solve for an action trajectory that satisfies the constraints on action derivatives while achieving greater smoothness, making it a trajectory optimization problem. $\mathbf{Q}$ also has a intuitive theoretical interpretation, which will be elaborated in detail in Section III-D. In Section III-D, we will dive deeper into the trajectory optimization objectives by incorporating learning-based perspectives.
The optimization problem (6) is a quadratic program. In DiffOG, we instantiate this class by proposing a specific, learnable trajectory optimization formulation tailored for imitation learning. Thus far, we have introduced the overall formulation of trajectory optimization in DiffOG. To further enable learning within this optimization layer, it is essential to emphasize a key property: differentiability. A differentiable trajectory optimization module can be seamlessly integrated with neural networks for end-to-end training, enabling the optimization process to acquire stronger representational capacity through data-driven learning. In what follows, we dive into the details of differentiability, beginning with the following proposition:
@@
[…]
@@
In terms of differentiability, by Proposition 1 and the work in [1], we can get the following remark.
@@
Since $\mathbf{Q}\succ 0$, the optimization problem (6) is strictly convex, ensuring a unique optimal solution $\mathbf{y}^{\star}_{t}$ that is continuous and subdifferentiable with respect to all variables everywhere, and differentiable at all but a measure-zero set of points. In this case, for example, the gradient $\nabla_{\mathbf{Q}}\mathbf{y}^{\star}_{t}$ can be computed explicitly via the solution of the KKT conditions, leveraging the fact that the optimality conditions are differentiable [1].
@@ III-D The Learning Objectives of DiffOG
First, let us assume that $\mathbf{Q}$ is a symmetric positive definite matrix. According to Proposition 1, this condition with proper constraints values ensures that the optimization problem (6) is both feasible and differentiable. By Cholesky factorization, there exists a unique upper-triangular matrix $\mathbf{R}$ with positive diagonal entries such that $\mathbf{Q}=\mathbf{R}^{\mathrm{T}}\mathbf{R}$. Because $\mathbf{Q}$ is symmetric positive definite, the factor $\mathbf{R}$ is invertible (i.e., full rank).
@@
which represent a linear transformation of the original action $\mathbf{a}_{t}$. In the presence of constraints and smoothing terms, the optimization problem (8) is no longer solely about a linear transformation; instead, it simultaneously accounts for three aspects: transformation, constraints, and smoothness. Next, we demonstrate how differentiability and learning can be leveraged to integrate these three aspects into three interpretable learning objectives.
@@
In this case, 1) the forward pass of DiffOG involves solving the optimization problem (8) to generate $\mathbf{y}^{\star}_{t}$, and 2) the backpropagation process updates $\mathbf{Q}$ using the loss function (10). The interpretable learning objectives can then be summarized in the following remark.
@@ III-E Parameterize $\mathbf{Q}$ with Transformer
By Remark 2, we know that optimization problem (6) is differentiable, and $\nabla_{\mathbf{Q}}\mathbf{y}^{\star}_{t}$ can be computed. The operations in Algorithm 1 are also differentiable. Then $\nabla_{\theta}{\mathbf{Q}}$ can be computed. Therefore, based on the chain rule, DiffOG can be fully trained in an end-to-end manner. For quadratic programming solver, we utilize qpth [1], which enables batch-form quadratic programming computation along with efficient backpropagation.
@@
    $\mathbf{L}\leftarrow\mathrm{tril}(\mathbf{L})$\ set upper-triangular elements to zero
@@
    $\mathbf{Q}\leftarrow\frac{1}{2}(\mathbf{Q}+\mathbf{Q}^{\mathrm{T}})$\ enhance symmetry
@@ IV Experiments
We validated DiffOG on 13 different tasks as shown in Fig. 3. Experiments involve two base policies, Diffusion Policy and 3D Diffusion Policy (DP3), along with a comprehensive comparison of trajectory processing baselines, including constraint clipping and penalty-based optimization. Additionally, we compare DiffOG with the constrained visuomotor policy [33]. We also conduct ablation studies on DiffOG.
For the hyperparameters of Diffusion Policy and DP3, we use the default parameters provided in their respective work [5, 35]. DiffOG, penalty-based optimization, and constraint clipping use same constraints and same base policy, either DP3 or Diffusion Policy, depending on the type of camera input used for the task. For details on the parameters of DiffOG and trajectory processing baselines, as well as information on the datasets, please refer to Appendix.
@@
Comparison with a Constrained Visuomotor Policy: We present comparative experiments between DiffOG and an constrained visuomotor policy Leto [33], showing that DiffOG achieves superior performance.
@@
[…]
TABLE I: Success rate of simulated tasks using Diffusion Policy [5] as the base policy. We report the mean success rate, calculated as the average over the final 10 checkpoints, with each checkpoint evaluated across 50 different environment initializations. The results are further averaged over 3 training seeds, totaling 150 environment evaluations. We use proficient-human demonstrations of Robomimic benchmark [13]. We highlight in bold the method that achieves the highest success rate after applying trajectory processing. DiffOG, penalty-based optimization, and constraint clipping use same constraints and same base policy. On challenging tasks such including square, tool hang, and transport, DiffOG outperforms penalty-based optimization with an average success rate improvement of 9% and surpasses constraint clipping with an average success rate improvement of 14.7%.
@@
Base Policy: DP3 0.98±0.010.98$\pm$ 0.010.98 ± 0.01 0.77±0.040.77$\pm$ 0.040.77 ± 0.04 0.87±0.020.87$\pm$ 0.020.87 ± 0.02 1.00±0.001.00$\pm$ $0$ 0.00 0.70±0.020.70$\pm$ 0.020.70 ± 0.02
[…]
TABLE II: Success rate of simulated tasks using DP3 [35] as the base policy. We report the mean success rate, calculated as the average over the final 10 checkpoints, with each checkpoint evaluated across 50 different environment initializations. The results are further averaged over 3 training seeds, totaling 150 environment evaluations. We highlight in bold the method that achieves the highest success rate after applying trajectory processing. All tasks are trained using 30 demonstrations. DiffOG, penalty-based optimization, and constraint clipping use same constraints and same base policy.
@@
TABLE III: Success rate of real tasks using Diffusion Policy [35] as the base policy and using ALOHA [36] as experimental platform. We highlight in bold the method that achieves the highest success rate.
@@
TABLE IV: We report the average inference time over 50 inferences for the base policy and DiffOG on 3080. For diffusion policy, we use 16 DDIM [23] steps for arrange desk and 8 DDIM steps for move the stack, with arrange desk requiring more due to its longer horizon and higher DoF. Results demonstrate DiffOG’s real-time feasibility.
@@
Figure 7: Comparison of trajectory metrics across methods on Robomimic and Meta-World benchmarks. These results are derived from 1500 rollouts (across 50 environments, 3 seeds, and the last 10 checkpoints per seed), providing strong statistical significance. Note that these values do not represent simple variations around a common mean and therefore cannot be interpreted in the same way as conventional mean $\pm$ standard deviation error bars, as they are not computed from a single rollout trajectory. We report them separately to highlight distinct aspects of trajectory quality. For Robomimic tasks, the larger maximum acceleration and greater standard deviation of acceleration over time indicate less smooth trajectories. Similarly, for Meta-World tasks, this observation applies to the velocity metrics. Since the actions provided by the Robomimic dataset are already normalized [13], the acceleration metrics shown here for Robomimic are also normalized. For details on the calculation of acceleration and velocity metrics, as well as the details to unnormalize the Robomimic metrics, please refer to Appendix.

### v5

@@ I Introduction
Imitation learning [pomerleau1988alvinn] has emerged as a popular paradigm for endowing robots with complex manipulation skills by leveraging human demonstrations. By formulating a supervised learning problem that maps sensor observations directly to actions via neural networks, imitation learning-based visuomotor policies have demonstrated effectiveness across a broad spectrum of tasks [chi2023diffusionpolicy, chi2024universal, zhao2023learning, fu2024mobile, shafiullah2023bringing]. Yet, these learned policies often exhibit notable limitations when applied to real-world settings, where safety, robustness, and strict constraint satisfaction are paramount. In contrast to classical model-based trajectory optimization methods which explicitly account for motion constraints and can ensure smooth and reliable trajectories [schulman2014motion], imitation learning approaches may produce suboptimal or jerky actions. Such trajectories can reduce the robustness of the whole robotic systems and raise critical safety concerns.
Meanwhile, various challenges in imitation learning, such as the representation of multi-modal action distributions [chi2023diffusionpolicy, lee2024behavior] and training large, generalized policies on extensive datasets [octo_2023, o2023open, black2024pi_0], have been the focus of ongoing research. Despite these advancements, ensuring the generation of high-quality and constraint-compliant trajectories for neural network-based policies remains a significant and open challenge. Unlike traditional model-based motion generation methods, learning-based approaches tend to be less controlled and more difficult to interpret. For instance, learning-based policies lack mechanisms to directly control trajectory smoothness or impose motion constraints through tunable parameters.
@@
We evaluated DiffOG across 11 simulated tasks and 2 real-world tasks. The results demonstrate that DiffOG significantly enhances the trajectory quality of visuomotor policies with generalizability while having minimal impact on policy performance, outperforming trajectory processing baselines such as constraint clipping and penalty-based trajectory optimization. Furthermore, DiffOG achieves superior performance compared to existing constrained visuomotor policy [xu2024leto].
@@ II-A Supervised Policy Learning
Supervised policy learning entails training a policy using supervised learning techniques on a dataset of pre-collected demonstrations [pomerleau1988alvinn]. Recent advancements in machine learning algorithms and data collection frameworks have substantially advanced the capabilities and effectiveness of supervised policy learning in manipulation applications. For example, various data collection frameworks and devices have been proposed in [chi2024universal, zhao2023learning, fu2024mobile, shafiullah2023bringing, seo2024legato, wang2024dexcap], enabling efficient data collection and the training of effective manipulation policies. To better capture the multi-modal action distributions present in demonstration data, prior work have explored various techniques, such as using diffusion models [chi2023diffusionpolicy], VQ-VAE [lee2024behavior], and energy-based models [florence2022implicit]. Furthermore, various effective action representations for supervised policy learning have been proposed, such as flow by point tracking [xu2024flow, wen2023any] and equivariant action representation [wangequivariant]. Moreover, many work have contributed to training more general and scalable policies on large-scale datasets [octo_2023, o2023open, black2024pi_0, liu2024rdt].
@@ II-B Policies with Differentiable Optimization
The integration of differentiable optimization into policy learning has garnered significant attention due to its potential to combine model-based formulations with the representational power of neural networks. Previous studies have investigated the integration of model-based structures with policies through differentiable optimization, including differentiable model predictive control (MPC) [amos2018differentiable] and the Koopman operator framework [retchin2023koopman]. However, these methods are predominantly designed for low-dimensional observations, which limits their flexibility and applicability to the general formulation of visuomotor policies. The work in [xu2024letac] presents a tactile-reactive grasping controller that integrates an image encoder with differentiable MPC. Although effective for specific tactile-based grasping tasks, this approach lacks the formulation required for broader policy learning frameworks. In the domain of obstacle avoidance, an end-to-end learning framework incorporating differentiable optimization is proposed in [xiao2021barriernet]. However, this approach is specifically designed for navigation tasks and lacks applicability to robotic manipulation scenarios and visuomotor policies.
@@
Riemannian motion policies [ratliff2018Riemannian] and their extensions [li2021rmp2, cheng2020rmp, rana2021towards] advance the generation of motions in high-dimensional spaces with complex, nonlinear dynamics into end-to-end robot learning by leveraging automatic differentiation and Riemannian geometry. However, in the context of policy learning, Riemannian motion policies are primarily designed for tasks involving low-dimensional observations, such as those demonstrated through kinesthetic teaching [rana2021towards]. In contrast, DiffOG is tailored for a broader range of visuomotor policies, refining policies that utilize camera observations as input. DiffOG focuses on generating task-completing trajectories while enhancing smoothness and ensuring compliance with hard constraints.
@@
DiffTORI [wan2024difftori] integrates differentiable trajectory optimization into the policy representation to generate actions, with a focus on capturing multi-modal action distributions in the context of imitation learning. However, DiffTORI lacks explicit and rigorous formulations for constraints and smoothness, limiting its ability to optimize trajectory smoothness and enforce compliance with hard constraints. In contrast, DiffOG incorporates a rigorously designed optimization layer that not only enables actions to satisfy hard constraints, but also enhances the interpretability of the policy.
@@
Leto [xu2024leto], by employing differentiable optimization, enables end-to-end training and inference that embedded trajectory optimization, allowing constraints to be incorporated as part of the training objectives. However, Leto was designed as a discriminative model, which limits its ability to represent multi-modal action distributions [florence2022implicit]. Its use of an optimization layer as the action head makes it difficult to integrate with commonly-used diffusion-based action heads [chi2023diffusionpolicy, octo_2023, liu2024rdt, ze20243d], significantly reducing its flexibility and applicability. In comparison, DiffOG offers greater flexibility and enhanced representational capacity, enabling it to achieve superior policy performance and adapt more effectively to a wider range of tasks.
@@ III-B Action Space
Actions fall into two categories: discrete actions, such as controlling grasping and releasing using discrete values [mandlekar2022matters], and continuous actions, such as the motion of the robot. Discrete actions, by their nature, cannot form a continuous trajectory and therefore do not require trajectory optimization. To address this distinction, we introduce a selection matrix $S\in\mathbb{R}^{D_{c}\times D_{a}}$ that identifies the action dimensions to be considered for trajectory optimization. Specifically, $S$ is used to select: 1) degrees of freedoms (DOFs) that are inherently continuous and capable of forming a trajectory, and 2) DOFs that require trajectory optimization in practical applications. The resulting filtered action variable is expressed as $c=Sa\in\mathbb{R}^{D_{c}}$. In this paper, we filter out the grasping action using $S$, as grasping typically does not require trajectory motion constraints and optimization.
Recently, considerable papers demonstrate that using the training objective of predicting a sequence of actions results in a more effective policy compared to predicting a single action [chi2023diffusionpolicy, zhao2023learning, lee2024behavior]. Therefore, at time step $t$, the action sequence output of policies adhere to the format of
@@ III-C Differentiable Trajectory Optimization
In optimization problem (6), $\mathbf{Q}$ is a weight matrix, $\alpha$ is a scalar, and $d_{\text{max}},d_{\text{min}}$ are constraint bounds for the derivatives of the robot actions. Intuitively, $\alpha$ reflects the degree to which the trajectory smoothness is emphasized. When the sum of the derivatives of the robot actions $\hat{\mathbf{c}}_{t}$ in the generated trajectory $\mathbf{y}_{t}$ is large, it indicates that the trajectory is less smooth. For instance, excessive acceleration or velocity contributes to an increase in the cost function. Therefore, we call $\alpha$ as smoothing weight. Optimization problem (6) aims to solve for an action trajectory that satisfies the constraints on action derivatives while achieving greater smoothness, making it a trajectory optimization problem. $\mathbf{Q}$ also has a intuitive theoretical interpretation, which will be elaborated in detail in Section III-D. In Section III-D, we will dive deeper into the trajectory optimization objectives by incorporating learning-based perspectives.
Optimization problem (6) is a quadratic program. In DiffOG, we instantiate this class by proposing a specific, learnable trajectory optimization formulation tailored for imitation learning. Thus far, we have introduced the overall formulation of trajectory optimization in DiffOG. To further enable learning within this optimization layer, it is essential to emphasize a key property: differentiability. A differentiable trajectory optimization module can be seamlessly integrated with neural networks for end-to-end training, enabling the optimization process to acquire stronger representational capacity through data-driven learning. In what follows, we dive into the details of differentiability, beginning with the following proposition:
@@
Because rank($S$) = $D_{c}$, rank($\mathbf{S}$) = $T_{p}D_{c}$. Because $c$ is taken from a subset of $a$, it follows that $D_{a}\geq D_{c}$. Therefore, $\mathbf{S}$ has full rank. Since $\mathbf{Q}$ is symmetric positive definite and the smoothing weight $\alpha\geq 0$, the matrix $\mathbf{Q}+\alpha\mathbf{S}^{\mathrm{T}}\mathbf{A}_{\text{diff}}^{\mathrm{T}}\mathbf{A}_{\text{diff}}\mathbf{S}$ is also symmetric positive definite. The condition $d_{\text{min}}<d_{\text{max}}$ ensures the feasibility of constraints, guaranteeing constraints are not contradictory.
@@
In terms of differentiability, by Proposition 1 and the work in [amos2017optnet], we can get the following remark.
@@
Since $\mathbf{Q}\succ 0$, optimization problem (6) is strictly convex, ensuring a unique optimal solution $\mathbf{y}^{\star}_{t}$ that is continuous and subdifferentiable with respect to all variables everywhere, and differentiable at all but a measure-zero set of points. In this case, for example, the gradient $\nabla_{\mathbf{Q}}\mathbf{y}^{\star}_{t}$ can be computed explicitly via the solution of the KKT conditions, leveraging the fact that the optimality conditions are differentiable [amos2017optnet].
@@ III-D The Learning Objectives of DiffOG
First, let us assume that $\mathbf{Q}$ is a symmetric positive definite matrix. According to Proposition 1, this condition with proper constraints values ensures that optimization problem (6) is both feasible and differentiable. By Cholesky factorization, there exists a unique upper-triangular matrix $\mathbf{R}$ with positive diagonal entries such that $\mathbf{Q}=\mathbf{R}^{\mathrm{T}}\mathbf{R}$. Because $\mathbf{Q}$ is symmetric positive definite, the factor $\mathbf{R}$ is invertible (i.e., full rank).
@@
which represent a linear transformation of the original action $\mathbf{a}_{t}$. In the presence of constraints and smoothing terms, optimization problem (8) is no longer solely about a linear transformation; instead, it simultaneously accounts for three aspects: transformation, constraints, and smoothness. Next, we demonstrate how differentiability and learning can be leveraged to integrate these three aspects into three interpretable learning objectives.
@@
In this case, 1) the forward pass of DiffOG involves solving optimization problem (8) to generate $\mathbf{y}^{\star}_{t}$, and 2) the backpropagation process updates $\mathbf{Q}$ using the loss function (10).
In both the DiffOG dataset training and DiffOG refine training, the differentiability of optimization problem (6) allows gradients from the loss to be propagated to $\mathbf{Q}$, as shown in Fig. LABEL:fig:firstPage. This enables the update of the optimization problem itself, making the inference process progressively adapt to the demonstration data. The interpretable learning objectives can then be summarized in the following remark.
@@ III-E Parameterize $\mathbf{Q}$ with Transformer
By Remark 2, we know that optimization problem (6) is differentiable, and $\nabla_{\mathbf{Q}}\mathbf{y}^{\star}_{t}$ can be computed. The operations in Algorithm 1 are also differentiable. Then $\nabla_{\theta}{\mathbf{Q}}$ can be computed. Therefore, based on the chain rule, DiffOG can be fully trained in an end-to-end manner. For quadratic programming solver, we utilize qpth [amos2017optnet], which enables batch-form quadratic programming computation along with efficient backpropagation.
@@
    $\mathbf{L}\leftarrow\mathrm{tril}(\mathbf{L})$ \\ set upper-triangular elements to zero
@@
    $\mathbf{Q}\leftarrow\frac{1}{2}(\mathbf{Q}+\mathbf{Q}^{\mathrm{T}})$ \\ enhance symmetry
@@ IV Experiments
We validated DiffOG on 13 different tasks as shown in Fig. 3. Experiments involve two base policies, Diffusion Policy and 3D Diffusion Policy (DP3), along with a comprehensive comparison of trajectory processing baselines, including constraint clipping and penalty-based optimization. Additionally, we compare DiffOG with the constrained visuomotor policy [xu2024leto]. We also conduct ablation studies on DiffOG.
For the hyperparameters of Diffusion Policy and DP3, we use the default parameters provided in their respective work [chi2023diffusionpolicy, ze20243d]. DiffOG, penalty-based optimization, and constraint clipping use same constraints and same base policy, either DP3 or Diffusion Policy, depending on the type of camera input used for the task. For details on the parameters of DiffOG and trajectory processing baselines, as well as information on the datasets, please refer to Appendix.
@@
Comparison with a Constrained Visuomotor Policy: We present comparative experiments between DiffOG and an constrained visuomotor policy Leto [xu2024leto], showing that DiffOG achieves superior performance.
@@
Base Policy: Diffusion Policy ${1.00}\pm{0.00}$ ${0.98}\pm{0.01}$ ${0.91}\pm{0.01}$ $0.83\pm 0.03$ $0.91\pm 0.01$ $0.84\pm 0.03$
DiffOG Dataset (Ours) $\textbf{1.00}\pm\textbf{0.00}$ $\textbf{0.98}\pm\textbf{0.01}$ $0.87\pm 0.02$ $\textbf{0.82}\pm\textbf{0.02}$ $0.89\pm 0.02$ $\textbf{0.83}\pm\textbf{0.02}$
DiffOG Refine (Ours) $\textbf{1.00}\pm\textbf{0.00}$ $\textbf{0.98}\pm\textbf{0.01}$ $\textbf{0.90}\pm\textbf{0.01}$ $0.81\pm 0.02$ $\textbf{0.91}\pm\textbf{0.03}$ $0.80\pm 0.02$
Constraint Clipping $\textbf{1.00}\pm\textbf{0.00}$ $0.93\pm 0.04$ $0.78\pm 0.03$ $0.59\pm 0.03$ $0.82\pm 0.02$ $0.82\pm 0.02$
Penalty-Based Optimization $\textbf{1.00}\pm\textbf{0.00}$ $0.96\pm 0.02$ $0.76\pm 0.03$ $0.79\pm 0.01$ $0.81\pm 0.02$ $0.81\pm 0.04$
TABLE I: Success rate of simulated tasks using Diffusion Policy [chi2023diffusionpolicy] as the base policy. We report the mean success rate, calculated as the average over the final 10 checkpoints, with each checkpoint evaluated across 50 different environment initializations. The results are further averaged over 3 training seeds, totaling 150 environment evaluations. We use proficient-human demonstrations of Robomimic benchmark [mandlekar2022matters]. We highlight in bold the method that achieves the highest success rate after applying trajectory processing. DiffOG, penalty-based optimization, and constraint clipping use same constraints and same base policy. On challenging tasks such including square, tool hang, and transport, DiffOG outperforms penalty-based optimization with an average success rate improvement of 9% and surpasses constraint clipping with an average success rate improvement of 14.7%.
@@
Base Policy: DP3 $0.98\pm 0.01$ $0.77\pm 0.04$ $0.87\pm 0.02$ $1.00\pm 0.00$ $0.70\pm 0.02$
DiffOG Dataset (Ours) $\textbf{0.98}\pm\textbf{0.01}$ $\textbf{0.73}\pm\textbf{0.01}$ $0.89\pm 0.05$ $\textbf{1.00}\pm\textbf{0.00}$ $\textbf{0.70}\pm\textbf{0.03}$
DiffOG Refine (Ours) $\textbf{0.98}\pm\textbf{0.01}$ $0.72\pm 0.02$ $\textbf{0.90}\pm\textbf{0.03}$ $\textbf{1.00}\pm\textbf{0.00}$ $0.68\pm 0.02$
Constraint Clipping $\textbf{0.98}\pm\textbf{0.01}$ $0.70\pm 0.01$ $0.86\pm 0.05$ $\textbf{1.00}\pm\textbf{0.00}$ $0.69\pm 0.02$
Penalty-Based Optimization $\textbf{0.98}\pm\textbf{0.01}$ $0.67\pm 0.03$ $0.86\pm 0.06$ $\textbf{1.00}\pm\textbf{0.00}$ $\textbf{0.70}\pm\textbf{0.03}$
TABLE II: Success rate of simulated tasks using DP3 [ze20243d] as the base policy. We report the mean success rate, calculated as the average over the final 10 checkpoints, with each checkpoint evaluated across 50 different environment initializations. The results are further averaged over 3 training seeds, totaling 150 environment evaluations. We highlight in bold the method that achieves the highest success rate after applying trajectory processing. All tasks are trained using 30 demonstrations. DiffOG, penalty-based optimization, and constraint clipping use same constraints and same base policy.
@@
TABLE III: Success rate of real tasks using Diffusion Policy [ze20243d] as the base policy and using ALOHA [zhao2023learning] as experimental platform. We highlight in bold the method that achieves the highest success rate.
@@
TABLE IV: We report the average inference time over 50 inferences for the base policy and DiffOG on 3080. For diffusion policy, we use 16 DDIM [song2021denoising] steps for arrange desk and 8 DDIM steps for move the stack, with arrange desk requiring more due to its longer horizon and higher DoF. Results demonstrate DiffOG’s real-time feasibility.
@@
Figure 7: Comparison of trajectory metrics across methods on Robomimic and Meta-World benchmarks. These results are derived from 1500 rollouts (across 50 environments, 3 seeds, and the last 10 checkpoints per seed), providing strong statistical significance. Note that these values do not represent simple variations around a common mean and therefore cannot be interpreted in the same way as conventional mean $\pm$ standard deviation error bars, as they are not computed from a single rollout trajectory. We report them separately to highlight distinct aspects of trajectory quality. For Robomimic tasks, the larger maximum acceleration and greater standard deviation of acceleration over time indicate less smooth trajectories. Similarly, for Meta-World tasks, this observation applies to the velocity metrics. Since the actions provided by the Robomimic dataset are already normalized [mandlekar2022matters], the acceleration metrics shown here for Robomimic are also normalized. For details on the calculation of acceleration and velocity metrics, as well as the details to unnormalize the Robomimic metrics, please refer to Appendix.